OpenStudy (anonymous):

solve for x.... 3(x-4)=-4x +30

OpenStudy (lilg132):

3x - 12 = - 4x + 30 3x + 4x = 30 + 12 7x = 42 7x / 7 = 42 / 7 x = 6 remember to click good answer

OpenStudy (anonymous):

thanx

OpenStudy (lilg132):

your welcome